# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 6.3. Educational Takeaway
Our work underscores the importance of incorporating ethics education for all roles within AI development teams. Educational programs should move beyond traditional technical training to include modules on ethical risk assessment, transparency, communication, and the practical implementation of governance frameworks. Building on prior studies that highlight the lack of formal training in ethical AI practices (Griffin et al., 2024; Agbese et al., 2023), our findings suggest that curriculum designers and organizational leaders should develop role-specific training programs that integrate both technical measures (e.g., bias audits and explainable AI techniques) and collaborative, interdisciplinary approaches to ethical decision-making. Given the evolving nature of AI ethics, it is essential to equip AI team members with educational modules that promote lifelong learning. As Balebako et al. (Balebako and Cranor, 2014) suggest, with appropriate guidance and well-designed educational tools, AI practitioners will be better prepared to navigate the complex trade-offs among innovation, trust, and compliance.
7. Conclusion
@@
This paper presented a large-scale mixed-method study examining the perceptions, practices, and risk mitigation strategies related to AI ethics across diverse roles in AI development teams. Our findings reveal that explainability, transparency, and role-specific familiarity with ethical guidelines foster trust and accountability in AI systems and promote the consideration and confidence of ethical principles in development. We observed differences in how various groups - from AM and RA to AD, QA, and ISec teams - understand and implement AI ethics principles. These disparities, which are further influenced by organizational size, geographic region, and gender, display the need for education and improved interdisciplinary collaboration towards ethical AI. While our study advances current understanding by highlighting both technical and organizational challenges, it also points to several avenues for future work, including longitudinal analyses, deeper investigations into the regulatory impacts on AI practices, and the development of frameworks for collaboration between various demographics that make up AI development teams.
8. Acknowledgments
@@
This research was supported by NSF Award #2238047 and a Maine Business School Research Award. We would like to thank Ersilda Cako, an undergraduate student at the University of Maine, for her valuable assistance in developing the survey questions and conducting the open coding analysis of the qualitative data.
Appendix A Definitions and Codebook
Table 21. Definition of AI Ethics Principles
@@ Appendix A Definitions and Codebook
Table 22. Taxonomy of Human-AI Activities
@@
Table 23. AI Ethics Principles Risk and Mitigation Code Book

### v2

@@ 5.3 Educational Takeaways
Our work underscores the importance of incorporating ethics education for all roles within AI development teams. Educational programs should move beyond traditional technical training to include modules on ethical risk assessment, transparency, communication, and the practical implementation of governance frameworks. Building on prior studies that highlight the lack of formal training in ethical AI practices [griffin2024ethical, agbese2023implementing], our findings suggest that curriculum designers and organizational leaders should develop role-specific training programs that integrate both technical measures (e.g., bias audits and explainable AI techniques) and collaborative, interdisciplinary approaches to ethical decision-making. Given the evolving nature of AI ethics, it is essential to equip AI team members with educational modules that promote lifelong learning. As Balebako et al. [balebako2014improving] suggest, with appropriate guidance and well-designed educational tools, AI practitioners will be better prepared to navigate the complex trade-offs among innovation, trust, and compliance.
5.4 Limitations
Similar to any survey study, our analysis relies on participants’ self-reported data. Thus, it may be affected by self-report bias, recall bias, and/or social desirability bias. During the consent process, we informed the participants that the survey pertained to AI practices, but we intentionally did not mention ethics to minimize the impact of self-selection biases. The Prolific user base may not fully represent the diverse population of AI development teams; thus, the survey may face recruitment bias. We diversified our sample by including participants from various AI-related forums, as well as those from Prolific. We used multiple screening questions to ensure that recruited participants have experience in AI development activities (Section 3.1.2). We adopted a conservative process to remove participants who appeared to provide AI-generated answers by carefully scrutinizing the write-in responses and investigating semantic qualities of LLM vs. human responses [tang2024science, munoz2024contrasting, guo2023how]. However, AI-generated responses may have influenced the study’s outcomes, and we may have excluded some professionals. We carefully framed our questions to avoid prompting biased responses and began the survey with general questions about AI practices to minimize the framing bias. Although we aimed to minimize the risks associated with answering follow-up write-in questions, some biases may still exist in those questions. We employed statistical analyses to ensure the broad applicability of our findings. To control for Type I errors in the presence of multiple hypothesis tests, we reported our results after employing the Bonferroni correction. Finally, we addressed standard threats to validity in empirical software engineering research. To mitigate construct validity threats, such as variable definitions that may be interpreted differently across demographics, and social desirability bias, we employed validated taxonomies (e.g., NIST, AI ethics principles defined in [worsdorfer2023aibill]) and avoided mentioning ethics during recruitment. To reduce internal validity threats, including confounding variables, we employed statistical methods such as mixed-effects models and LASSO regression to isolate predictive factors across demographics. External validity was strengthened by recruiting participants from multiple sources, including crowdsourcing platforms, and filtering for relevant AI development experience via Prolific. Finally, to enhance conclusion validity, we filtered out low-quality or AI-generated responses and applied Bonferroni correction to minimize Type I errors during hypothesis testing.
6 Conclusion
@@
This paper presented a large-scale mixed-method study examining the perceptions, practices, and risk mitigation strategies related to AI ethics across diverse roles in AI development teams. Our findings reveal that explainability, transparency, and role-specific familiarity with ethical guidelines foster trust and accountability in AI systems, promoting the consideration and confidence in ethical principles during development. We observed differences in how various groups, from AM and RA to AD, QA, and ISec teams, understand and implement AI ethics principles. These disparities, which are further influenced by organizational size, geographic region, and gender, underscore the need for education and enhanced interdisciplinary collaboration in the development of ethical AI. While our study advances the current understanding by highlighting both technical and organizational challenges, it also points to several avenues for future work, including longitudinal analyses, deeper investigations into the regulatory impacts on AI practices, and the development of frameworks for collaboration among the various demographics that comprise AI development teams.
[…]
Supplementary information
We have published a replication package, available at https://zenodo.org/records/17905482, which contains all responses and the code for any analysis presented in this paper.
[…]
Acknowledgements
@@
This research was supported by NSF Award #2238047 and a Maine Business School Research Award. We would like to thank Ersilda Cako, an undergraduate student at the University of Maine, for her valuable assistance in developing the survey and conducting the open coding analysis of the qualitative data.
Appendix A Participants Demographcis
Table 27: Breakdown of Participants’ Demographics
Age 18-25 (13.0%) 26-35 (50.5%) 36-45 (24.4%) 46+ (11.8%) PnS (0.3%)
Appendix B Definitions and Taxonomy
Table 28: Definition of AI Ethics Principles
@@ Appendix B Definitions and Taxonomy
Table 29: Taxonomy of Human-AI Activities
@@
Appendix C Codebook
Table 30: AI Ethics Principles Risk and Mitigation Codebook
@@ Appendix C Codebook
Appendix D LASSO Regression Predictors
Table 31: LASSO Regression Predictor Categories by Participant Group
Predictor Category Group A Group B
Demographics
     Age $\bullet$ $\bullet$
     Gender $\bullet$ $\bullet$
     Education Level $\bullet$ $\bullet$
     Degree Field $\bullet$ $\bullet$
     Company Size $\bullet$ $\bullet$
     Company Type $\bullet$ $\bullet$
     Location $\bullet$ $\bullet$
     Role $\bullet$ $\bullet$
     Current Job Experience $\bullet$ $\bullet$
     AI Development Experience $\bullet$
AI Practice
     Company Builds/Deploys AI $\bullet$ $\bullet$
     Currently Developing AI $\bullet$ $\bullet$
     AI Usage Frequency $\bullet$ $\bullet$
     AI Applications (multi-label) $\bullet^{a}$ $\bullet$
     AI Effects/Impacts (multi-label) $\bullet^{a}$ $\bullet$
     AI Tools/Technologies $\bullet$
     Ethics Consideration Frequency $\bullet$
     Mitigation Strategies $\bullet$
Perception
     Regulation Impact $\bullet$ $\bullet$
     Regulation Negative Impact $\bullet$ $\bullet$
     Principles Considered $\bullet$
     Principles at Risk $\bullet^{a}$ $\bullet$
Total Features 68–78b 141–142
Sample Size (N) 76–90 141–178
aIncluded in Track A but excluded in some models to prevent overfitting. bFeature count varies; multi-label questions are one-hot encoded.